DETERMINE THE RESISTANCE OF A BUSBAR MADE OF COPPER IF THE LENGTH IS 5m LONG


AND THE CROSS-SECTION IS A 4X4š‘m^2 . USE 1.7241 MICRO OHM-CM.

1
Expert's answer
2021-11-17T11:25:23-0500

We have the parameters:

Resistance(R)= ?

Resistivity (Ļ) =1.7241micro Ī©-cm

Area(A) =Ā "4*4cm\u00b2"

Length (L)=Ā "5m"

The formula of resistance R is given by:

"R=\frac{ρL}{A}"

We convert resistivity fromĀ micro Ī©-cmĀ toĀ micro Ī©-m

"\frac{1.724}{100}=0.01724micro Ω-m"

We convert "4*4cm\u00b2" to "m\u00b2"

"\frac{4*4}{10000}m²=0.0016m²"


Now, we substitute the parameters in the formulaĀ "R=\frac{ρL}{A}"

"R=\frac{0.017241*5}{0.0016}"


"R=\frac{0.086205}{0.0016}"


"R=53.878125\u03a9"

"R=53.88\u03a9(2.d.p)"